OpenStudy (anonymous):

How to integrate x^5sqrt(x^2+3)

OpenStudy (raden):

int (x^5sqrt(x^2+3)) = int ((x^2)^2*x*sqrt(x^2+3))dx let u = x^2+3 -----> x^2 = u-3 du = 2x dx 1/2 du = x dx continue it ....

OpenStudy (anonymous):

thanks

OpenStudy (raden):

yw